OpenStudy (anonymous):

Please help! What is the slope of the line passing through the points (-1, 7) and (-1, 3)?

OpenStudy (anonymous):

slope= (y2-y1)/(x2-x1)

OpenStudy (anonymous):

\[Slope = (y_2 - y_1)/(x_2 - x_1)\] Here, x1 = -1, y1 = 7 x2 = -1 and y2 = 3 Put these values and solve...

OpenStudy (anonymous):

it will be infinity because n/0 = infinity

OpenStudy (anonymous):

are you sure the problem was this ???

OpenStudy (anonymous):

REMEMBER: IF SLOPE COMES OUT TO BE INFINITY, THEN Slope is basically given as: \[Slope(m) = \tan \theta\] If m = Infinity, then\[\tan \theta = \infty\] Therefore, angle(theta) will be 90... As tan90 = Infinity...

OpenStudy (anonymous):

THANKS